Compute the dimensions of a rectangular flower garden with a perimeter of 250m and with length that exceeds three times its width by 5. pls help. :(
Answer by blwinbbbles(106)

OK start with what you know: Use l for length and w for width
perimeter is the outside distance around the rectangle
So:
Since there are 2 lengths and 2 widths on a rectangle then
2w + 2l = 250
and we know that length is 5 more than 3 times the width
so: l = 3w +5
Now use substitution:
2w + 2(3w + 5) = 250
solve:
2w + 6w + 10 = 250 subtract 10 from both sides and add the w's together
8w = 240 divide by 8
w = 30
so if w = 30 then length is l = 3W + 5
l = 3(30) + 5
l = 90 +5
l = 95
So your answer is width is 30m and length is 95m
Hope this helps